# DRP: Distilled Reasoning Pruning with Skill-aware Step Decomposition for Efficient Large Reasoning Models

### v1

@@ Abstract
While Large Reasoning Models (LRMs) have demonstrated remarkable success in complex reasoning tasks through Long Chain-of-Thought (CoT) reasoning, their inference often involves excessively verbose reasoning traces, resulting in substantial inefficiency. To address this issue, we propose Distilled Reasoning Pruning (DRP), a hybrid framework that combines inference-time pruning with tuning-based distillation—two widely used strategies for efficient reasoning. DRP employs a teacher model to perform skill-aware step decomposition and content pruning, and then distills the pruned reasoning paths into a student model, enabling it to reason both efficiently and accurately. Across a series of challenging mathematical reasoning datasets, we find models trained with DRP achieve substantial improvements in token efficiency without sacrificing accuracy. Specifically, DRP reduces the average token usage on GSM8K from 917 to 328 while improving accuracy from 91.7% to 94.1%, and achieves a 43% token reduction on AIME with no performance drop. Further analysis reveals that aligning the reasoning structure of training CoTs with the student’s reasoning capacity is critical for effective knowledge transfer and performance gains.
@@ 1 Introduction
Although Large Reasoning Models (LRMs) Xu et al. (2025a), like OpenAI’s o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), have advanced the state of the art in complex reasoning tasks Li et al. (2025b), a critical limitation of these models is their tendency toward overthinking—the generation of excessively verbose reasoning trajectories containing redundant or unnecessary steps Chen et al. (2024); Cui et al. (2025); Fu et al. (2024). This can lead to substantial inference overhead and misguide the model toward incorrect conclusions Sui et al. (2025).
Existing solutions predominantly follow two paradigms: inference-time pruning, which attempts to terminate generation early to avoid redundant reasoning steps Cui et al. (2025); Muennighoff et al. (2025); Fu et al. (2024); Zeng et al. (2025), and distillation-based compression, where smaller models are trained on teacher-generated reasoning paths to imitate the concise reasoning behavior of larger models Xu et al. (2025b); Zhu et al. (2024). However, both approaches can be at the cost of accuracy: pruning methods risk prematurely halting the reasoning process, thereby preventing the model from reaching the correct answer, while distillation methods tend to underperform—particularly for small-scale student models—due to the learnability gap, which hinders effective knowledge transfer when the performance or size gap between student and teacher is too great Xu et al. (2024); Li et al. (2025a).
@@ 2.1 The Overthinking Problem
Large reasoning models, such as OpenAI o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), are a specialized subclass of large language models trained to iteratively generate intermediate reasoning steps and refine the thinking process until reaching a final answer Chen et al. (2025); Sui et al. (2025). Unlike traditional LLMs, which rely on test-time strategies such as Chain-of-Thought prompting to elicit detailed reasoning, LRMs are trained to internalize CoT generation. This design significantly enhances their performance on reasoning-intensive tasks, particularly in challenging mathematics and programming. However, LRMs, especially those with smaller parameter sizes, often exhibit a tendency to produce overly verbose reasoning chains, including unnecessary tracebacks and redundant alternative paths Chen et al. (2024); Sui et al. (2025); Cui et al. (2025); Fu et al. (2025, 2024). Such overthinking not only increases token consumption beyond practical inference budgets but also introduces reasoning drift that can degrade output correctness Hou et al. (2025); Sui et al. (2025). To address these, our method explicitly prunes harmful overthinking behaviors while simultaneously guiding the model toward concise and effective reasoning expressions.
@@ 4.5 Evaluation Protocol
We use the lmms-eval [^2] https://github.com/EvolvingLMMs-Lab/lmms-eval framework for unified evaluation across tasks. Each model is evaluated in a zero-shot setting. For inference, we use the vLLM backend with the maximum generation length set to 131,072 tokens—the upper limit supported by the Qwen models.
@@ Appendix D Skill Base Decomposition Evaluation
To assess the quality of our skill-based step decomposition, we conduct a pairwise comparison experiment using Gemini 2.0 Flash as the judge. For each reasoning trace $T$, we generate two segmentations: (1) our proposed skill-based decomposition, and (2) a baseline sentence-based split (obtained by naïvely splitting at punctuation or sentence boundaries).

### v2

@@ Abstract
While Large Reasoning Models (LRMs) have demonstrated remarkable success in complex reasoning tasks through Long Chain-of-Thought (CoT) reasoning, their inference often involves excessively verbose reasoning traces, resulting in substantial inefficiency. To address this issue, we propose Distilled Reasoning Pruning (DRP), a hybrid framework that combines inference-time pruning with tuning-based distillation—two widely used strategies for efficient reasoning. DRP employs a teacher model to perform skill-aware step decomposition and content pruning, and then distills the pruned reasoning paths into a student model, enabling it to reason both efficiently and accurately. Across a series of challenging mathematical reasoning datasets, we find models trained with DRP achieve substantial improvements in token efficiency without sacrificing accuracy. Specifically, DRP reduces the average token usage on GSM8K from 917 to 328 while improving accuracy from 91.7% to 94.1%, and achieves a 43% token reduction on AIME with no performance drop. Further analysis reveals that aligning the reasoning structure of training CoTs with the student’s reasoning capacity is critical for effective knowledge transfer and performance gains. We release our code at: https://github.com/YuxuanJiang1/DRP
@@ 1 Introduction
Although Large Reasoning Models (LRMs) Xu et al. (2025a), like OpenAI’s o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), have advanced the state of the art in complex reasoning tasks Li et al. (2025c), a critical limitation of these models is their tendency toward overthinking—the generation of excessively verbose reasoning trajectories containing redundant or unnecessary steps Chen et al. (2024); Cui et al. (2025); Fu et al. (2024). This can lead to substantial inference overhead and misguide the model toward incorrect conclusions Sui et al. (2025).
Existing solutions predominantly follow two paradigms: inference-time pruning, which attempts to terminate generation early to avoid redundant reasoning steps Cui et al. (2025); Muennighoff et al. (2025); Fu et al. (2024); Zeng et al. (2025), and distillation-based compression, where smaller models are trained on teacher-generated reasoning paths to imitate the concise reasoning behavior of larger models Tan et al. (2024); Zhu et al. (2024); Xu et al. (2025b). However, both approaches can be at the cost of accuracy: pruning methods risk prematurely halting the reasoning process, thereby preventing the model from reaching the correct answer, while distillation methods tend to underperform—particularly for small-scale student models—due to the learnability gap, which hinders effective knowledge transfer when the performance or size gap between student and teacher is too great Xu et al. (2024); Li et al. (2025b).
@@ 2.1 The Overthinking Problem
Large reasoning models, such as OpenAI o1 OpenAI (2024b) and DeepSeek-R1 Guo et al. (2025), are a specialized subclass of large language models trained to iteratively generate intermediate reasoning steps and refine the thinking process until reaching a final answer Chen et al. (2025); Sui et al. (2025); Yu et al. (2025). Unlike traditional LLMs, which rely on test-time strategies such as Chain-of-Thought prompting to elicit detailed reasoning Tong et al. (2024), LRMs are trained to internalize CoT generation. This design significantly enhances their performance on reasoning-intensive tasks, particularly in challenging mathematics and programming. However, LRMs, especially those with smaller parameter sizes, often exhibit a tendency to produce overly verbose reasoning chains, including unnecessary tracebacks and redundant alternative paths Chen et al. (2024); Sui et al. (2025); Cui et al. (2025); Fu et al. (2025, 2024). Such overthinking not only increases token consumption beyond practical inference budgets but also introduces reasoning drift that can degrade output correctness Hou et al. (2025); Sui et al. (2025). To address these, our method explicitly prunes harmful overthinking behaviors while simultaneously guiding the model toward concise and effective reasoning expressions.
@@ 4.5 Evaluation Protocol
We use the lm-evaluate-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness framework for unified evaluation across tasks. Each model is evaluated in a zero-shot setting. For inference, we use the vLLM backend with the maximum generation length set to 131,072 tokens—the upper limit supported by the Qwen models.
@@ Appendix D Skill Base Decomposition Evaluation
To assess the quality of our skill-based step decomposition, we conduct a pairwise comparison experiment using Gemini 2.0 Flash as the judge Gao et al. (2023); Li et al. (2024, 2025a). For each reasoning trace $T$, we generate two segmentations: (1) our proposed skill-based decomposition, and (2) a baseline sentence-based split (obtained by naïvely splitting at punctuation or sentence boundaries).